OpenStudy (belleuuuuuuuuu):

1. Sage correctly found the following quotient. 11/2 divided by 1/3= 4 1/2. (a) Create a model to show the quotient. (b) Explain what the half in the quotient refers to in your model. please help this is for my extra credit

OpenStudy (hoot_hoot):

\[\frac{ \frac{ 11 }{ 2 } }{ \frac{ 1 }{ 3 } }\ =\frac{ 11 }{ 2 } \div \frac{ 1 }{ 3 } = \frac{ 11 }{ 2 } \times 3 \neq 4\frac{ 1 }{ 2} \neq \frac{ 9 }{ 2 }\] Wait... Isn't this supposed to be in the Math section?